6. Suppose that an automobile manufacturer advertises that its new hybrid car has a mean gas mileage of 50 miles per gallon. You take a simple random sample of n = 30 hybrid vehicles and test their gas mileage. You find that in this sample, the average is ̄x = 47 miles per gallon with a standard deviation of s = 5.5 miles per gallon. Does this indicate that the advertiser’s statement is true or false? Let a=.0.05
1
Expert's answer
2019-11-14T09:44:12-0500

Let "X=" the gas mileage in miles per gallon

Assume that "\mu_0=50\ \text{miles per gallon}."

"H_0:\mu=\mu_0"

"H_1:\mu\not=\mu_0"

Given that "\sigma=5.5\ \text{miles per gallon}, n=30, \bar{X}=47\ \text{miles per gallon}"

From the Central Limit Theorem, we know this sampling distribution is normal with a mean of 50 and standard error of

"\sigma/\sqrt{n}=5.5/\sqrt{30}"

This corresponds to a two-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

Let "\alpha=0.05"

The critical value for a two-tailed test is "z_c=1.96"

The rejection region for this two-tailed test is "R=\{{z:|z|>1.96}\}"

The z-statistic is computed as follows:


"z={\bar{X}-\mu_0 \over \sigma/\sqrt{n}}"

"z={47-50 \over5.5/\sqrt{30}}\approx-2.9876"

Since it is observed that "|z|=|-2.9876|=2.9876>1.96=z_c," it is then concluded that the null hypothesis is rejected.

The p-value


"p=2\cdot P(Z<-2.9876)\approx2(0.0014)=0.0028"

Since "p=0.0028<0.05," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu" is different than "50," at the "0.05" significance level.

Confidence interval


"\bar{X}\pm z_c{\sigma \over \sqrt{n}}"

"47\pm 1.96{5.5 \over \sqrt{30}}"

"47\pm 1.968"

"(45.032,48.968)"